A resistance thermometer made from platinum has a resistance of R0=50 ohms at T0=20 degrees Celsius.(Temp coefficient of platinum, a=4.0 x 10^-3)
(a) When we connect it to a battery of 9V(no internal resistance), calculate the current.
(b)When the thermometer is immersed in hot water, we measure the current as 0.15A. hat is the temperature of the water?
1
Expert's answer
2019-06-24T09:30:24-0400

the current equals to

"I = \frac{E}{R} = \frac{9}{50} = 0.18 (A);"

the temperture is

"T = \frac{U-IR_0}{IR_0\alpha} + T_0 = \frac{9-7.5}{30}\cdot1000 + 20 = 50 + 20 =70 (C)"